eBook Find the interest rates earned on each of the following. Round your answers to the nearest whole number

. You borrow $700 and promise to pay back $742 at the end of 1 year

. % You lend $700 and the borrower promises to pay you $742 at the end of 1 year.

% You borrow $87,000 and promise to pay back $215,409 at the end of 8 years.

% You borrow $11,000 and promise to make payments of $3,359.50 at the end of each year for 5 years

. %
Answer #1
Future value = present value*(1+ rate)^time
742 = 700*(1+Interest rate/100)^1
Interest rate % = 6
Future value = present value*(1+ rate)^time
742 = 700*(1+Interest rate/100)^1
Interest rate % = 6
Future value = present value*(1+ rate)^time
215409 = 87000*(1+Interest rate/100)^8
Interest rate % = 12
PVOrdinary Annuity = C*[(1-(1+i/100)^(-n))/(i/100)]
C = Cash flow per period
i = interest rate
n = number of payments
11000= 3539.5*((1-(1+ Interest rate/100)^-5)/(Interest rate/100))
Interest rate% = 18.28